OpenStudy (anonymous):

convert the following into an exponential equation: log81^3=1/4

OpenStudy (lalaly):

base 8?

OpenStudy (anonymous):

i believe so

OpenStudy (lalaly):

\[\large{\log_8{1^3}}=\frac{1}{4}\]?

OpenStudy (anonymous):

no. log 81^3=1/4

OpenStudy (anonymous):

81 base

hero (hero):

\[\log_{81}3= ?\]

hero (hero):

http://www.wolframalpha.com/input/?i=log_ {81}%283%29

OpenStudy (anonymous):

3

OpenStudy (anonymous):

From the above information base can be worked out and it is 3^48

OpenStudy (anonymous):

the exponential notation of given equation is 3^12 = 81^3